Journalist Report – February 24th

Sol 12 – When stomachs are full, hearts are too

“Mars could do nothing to them, for they were bred to plains and prairies as open as the Martian fields. They came and made things a little less empty, so that others would find courage to follow.”

– Chapter 12 of The Martian Chronicles by Ray Bradbury

“HabCom, for the record, I have never seen a smile so wide on Alexandre’s face!”

I was sorting the first photos of the day sitting at the Hab table, when I overheard the radio communication between Alice, our HabCom, and Quentin, out on EVA to deploy the weather station, MegaAres and LOAC once and for all (fingers crossed!). MegaAres is finally fully functional! Up until now, Mars had not brought us much luck: MegaAres had to be retrieved for repairs, we had to remove the LOAC because of the rain, and the weather station was knocked out by the wind. Today, all instruments are fully deployed and operational! Even the field mill, an instrument designed to measure the electric field around the weather station, which was left uninstalled because of cabling issues, is up and running. Alexandre was ecstatic, and did a little dance around the antenna at the end of the EVA.

After lunch, the food and water resupply ships landed not far from the station, and moving between the different modules of the station was temporarily suspended. Corentin and I, stuck in the Science Dome for a little while, talked about how difficult it is to continue communicating with our loved ones, our lives being for the most part disconnected from Earth.

As my heart rate was being monitored for 30 very long minutes for an experiment, and as I stood there doing nothing, I realized once again how deformed our perception of time can be. Mid-rotation is just around the corner, time has gone by much too quickly for my taste, and here I am, 5 minutes into my ECG, feeling like I have been standing there for hours…

Upon returning to the Hab, I was so relieved to see cupboards and cabinets full of food! Watching Alice and Jérémy play catch with cereal boxes was a very endearing sight.

I feel like everyone is feeling a bit weary tonight: Corentin and Jérémy are working on solving a data handling issue, while Alice is feverishly preparing the schedule for the next week, and Quentin is trying to implement a location tracking system, essential to our experiment in partnership with the Swedish Royal Institute of Technology. Nevertheless, we maintain our optimist outlook on the mission and are glad that all atmospheric experiments are finally deployed. Hopefully, Alexandre will have both more time and good weather in the next few weeks to focus on his astronomy project!

To inaugurate our new, freshly off-the counter freeze-dried food stock, Jérémy is preparing a lasagna feast for the whole crew!

EVA Report – February 24th

Crew 275 EVA Report 24Feb2023

EVA # 11

Author: Alexandre VINAS (crew astronomer and EVA leader)

Purpose of EVA: Repairs of the weather station, installation of the antenna of the MegaAres instrument & Maintenance of the other atmospheric instruments

Start time: 9:00

End time: 11:45


9:02 Beginning of the depressurization

9:23 EVA crew arrived at destination

10:24 Installation of the antenna of MegaAres

10:37 All tests on the antenna are validated, the installation is a success

10:55 Purple Air and LOAC instrument turned on

11:11 An helicopter flew over us twice and over the MDRS at a altitude of approximately 50-70 meters

11:18 Field mill turned on

All the instruments are installed correctly, the EVA is a success.

11:45 EVA crew back to the main hab

Destination: Between the Hab and Marble Ritual near the road.

Coordinates (use UTM WGS 84): N4251000 E518500

Participants: EVA leader: Alexandre Vinas (Crew Astronomer), Marie Delaroche (Crew Journalist), Corentin Senaux (HSO)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual (see map on the attached picture)

Mode of travel: Driving, even if the destination is less than 1 km away, the crew will need to transport heavy instruments that also take up a lot of space, and then walking around the instruments.

GreenHab Report – February 24th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 72.2 °F

Average humidity: 31 %

Hours of supplemental light: 0 hour

Daily water usage for crops: 12.55 gallons

Daily water usage for research and/or other purposes: 0.13 gallons

Water in the blue tank: 52.22 gallons (before refill)

Time(s) of watering for crops: 9:54 & 13:58

Changes to crops: N/A

Narrative: Today’s morning I was on an EVA so it was the HSO that watered and maintained the plants and crops.

I, however, went to water them again in the afternoon.

Everything looks good and there is nothing to signal.

So far everything goes to plan, and crops are doing good. We could say, so far so good!

Harvest: N/A

Support/supplies needed: There is a thermal insulation on the GreenHab floor, I do not know if it belongs there. A crew member asked me if it is okay to use a little bit of it to insulate a battery for our scientific instrument or does it belong to the GreenHab for a Sun-related purpose?

Operations Report – February 24th

Crew 275 Operations Report 24Feb2023
SOL: 12

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• GreenHab Carbon Monoxide alarm

Notes on non-nominal systems:
• The GreenHab Carbon Monoxide whistled shortly three times in a row, indicating it has low battery. As we pushed the button to test it, the plastic button broke. The alarm is indeed out of battery.

Spirit rover used: No
• Hours: 209.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.4
• Beginning charge: 100%
• Ending charge: 94%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.2
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.8
• Beginning charge: 100%
• Ending charge: 95%
• Currently charging: Yes

General notes on rovers: Rovers were used during today’s EVA. Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. There was a short power outage at 7:45 am, quickly solved by Mission Support as soon as we advised. We received food resupply and water resupply. Following a misunderstanding, the Engineering Airlock was used today during the EVA, in order to transfer equipment outside, after a depressurization in this airlock. This event is described more in details in the Sol Summary.
• WATER USE: Impossible to say because I do not know the amount of water resupplied
• Water: 540.73 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Suits and radios were used during today’s EVA. Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations. Water was resupplied.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: No
• Harvest: None

Summary of ScienceDome operations: The Crew Scientist conducted routine operations, and tests for the Human Factors experiment were performed.
• Dual split: Switched off during the night

Summary of RAM operations: We performed thermal and humidity insolation of batteries.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: The GreenHab Carbon Monoxide whistled shortly three times in a row, indicating it has low battery. As we gently pushed the button to test it, the plastic button broke. The alarm is indeed out of battery.

Questions, concerns, and requests to Mission Support:
• Do you want us to put the GreenHab Carbon Monoxide in the RAM?
• Do you have any information about the amount of water resupplied in the Hab tank?

Sol Summary Report – February 24th

Crew 275 Sol Summary Report 24Feb2023

Sol: 12

Summary Title: When stomachs are full, hearts are too!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

The day started with a big misunderstanding regarding the conditions of the simulation. Indeed, even if we are very aware that a pressurized tunnel starts from the door of the engineering airlock, we were using it in the context of EVA to store material to be retrieved by EVA crews when there was not enough space in the main airlock for both the EVA crew and the material. We were very aware that this was a limitation of the simulation, but we were considering this as a virtual “third airlock” dedicated to transfer of material from the inside to the outside of the Hab, while the crew was either in the Hab or in a pressurized suit. I have been instructed to do so during my first mission at the MDRS in 2016, and last year’s Supaero crew was also operating this way. I reckon that the good practice would have been to include this in the EVA request, so that we could have had this important discussion beforehand. We have never tried to hide this procedure from Mission Support. For all our EVAs dedicated to the atmospheric instruments, our communications (in English) on channel 1 included sentences like “Request permission to open the door of the Engineering Airlock”, but I understand that this may have been lost in the middle of communications in French between the EVA crewmembers on channel 2. We now understand very well that we are not allowed to use the Engineering Airlock this way. From now on, do you confirm that we must wait for 5 minutes of re-pressurization to open the internal door of the main airlock, before putting the equipment there, and then wait another 5 minutes of depressurization before the EVA crew is allowed to open the external door of the main airlock and retrieve the equipment?

Besides this misunderstanding, the EVA was very successful: both the re-deployment of the instruments and their functional tests have been successful. We now have a fully functional set of atmospheric instruments.

The other good news of the day is the food resupply. It is always very impressive to observe how food can boost the crew’s morale!

Look Ahead Plan:

Tomorrow (Sol 13), we would like to perform our first EVA dedicated to the MetMet experiment. The plan would be to collect geological samples based on the information given by a user-friendly instrument measuring the magnetic susceptibility and conductivity of rocks (the aim being to check the interest of a rock sample before collecting it). In addition to this, we would change the batteries of the atmospheric instruments, because we do not plan to go on any EVA on Sunday (Sol 14), and the batteries cannot last more than 48 hours in cold weather.

Next week, we would like to plan 3 additional EVAs dedicated to photogrammetry: One for the mapping of the area, one for the exploration with a 2D map, and the third one for the exploration with a 3D render. This time the target area would be Candor Chasma.

Anomalies in work: None.

Weather: Windy and cloudy, low -3°C/27°F, high 5°C/41°F.

Crew Physical Status: Optimal.

EVA: EVA #11 was successful. All the objectives have been accomplished and we now have a fully operational set of atmospheric instruments.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site